OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 1 - 2i is a zero of f(x) = x4 - 2x3 + 6x2 - 2x + 5.

OpenStudy (anonymous):

you know how to find a quadratic polynomial with zeros \(1-2i\) and \(1+2i\) ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok there are a few ways one is easy, the other is really really easy lets do the easy way first

OpenStudy (anonymous):

working backwards from \[x=1+2i\] subtract 1 and get \[x-1=2i\] then square (carefully) \[(x-1)^2=(2i)^2\\ x^2-2x+1=-4\] then add \(4\) and get \[x^2-2x+5\]

OpenStudy (anonymous):

the real real easy way is to memorize that if \(a+bi\) is a zero of a quadratic, then it is \[x^2-2ax+(a^2+b^2)\] so in your case, since the zero is \(1+2i\) it is \[x^2-2\times 1+(1^2+2^2)=x^2-2x+5\]

OpenStudy (anonymous):

wow thankyou

OpenStudy (anonymous):

last job is to factor \[x^4 - 2x^3 + 6x^2 - 2x + 5=(x^2-2x+5)(something)\]

OpenStudy (anonymous):

i didnt look at it like that

OpenStudy (anonymous):

factoring should be pretty easy now right?

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

k good you should get \[(x^2-2x+5)(x^2+1)\] so other zeros are \(\pm i\)